How much water should be added to a 72 ounces of a 10% salt solution to make a 3% salt solution?
Answer by edjones(8007)

72*.1=7.2
x*0=0
(x+72)*.03=.03x+2.16
.03x+2.16=7.2+0
.03x=5.04
x=168 oz water.
.
Ed